# Discrete quantum groups and their duals

### v1

@@ Acknowledgments
The first, probably the main difference is that in [VD-discrete] the existence of the invariant integrals is proven from the axioms. In [Po-Wo], the existence of the integrals on the discrete quantum group is obtained given the fact that it is the dual of a compact quantum group. However, if you want to show that every discrete quantum group is the dual of a compact one, you need the integrals to do that.
@@
The second difference, of another nature but maybe equally important, is the fact that we consider the duality between compact and discrete quantum groups as a special case of the duality for algebraic quantum groups. In this theory, the general case turns out to be simpler than the special case. If e.g. the general construction of the dual (as for algebraic quantum groups) is applied to the special cases of a compact or a discrete quantum group, it is easy to see that the result is a discrete, respectively a compact quantum group. The arguments are simpler than the ones given in [Po-Wo].
This leads in general, also at other places where recently other aspects of compact and discrete quantum groups are studied, to simplifications and a better understanding. This should not come as a surprise because it is already the case in the approach to discrete quantum groups itself as treated in [Po-Wo].
@@
Taken abstraction of this (and of the historical context), people who want to work with compact and discrete quantum groups can consider any of the two approaches. To a great extend, it is a matter of taste which you take as a starting point (although we are convinced that starting with discrete quantum groups is a bit easier). However, for a full understanding, we feel it is worthwile to study the two theories also as independent entities.

### v2

@@ Acknowledgments
The first, probably the main difference is that in [28] the existence of the invariant integrals is proven from the axioms. In [18], the existence of the integrals on the discrete quantum group is obtained given the fact that it is the dual of a compact quantum group. However, if you want to show that every discrete quantum group is the dual of a compact one, you need the integrals to do that.
@@
The second difference, of another nature but maybe equally important, is the fact that we consider the duality between compact and discrete quantum groups as a special case of the duality for algebraic quantum groups. In this theory, the general case turns out to be simpler than the special case. If e.g. the general construction of the dual (as for algebraic quantum groups) is applied to the special cases of a compact or a discrete quantum group, it is easy to see that the result is a discrete, respectively a compact quantum group. The arguments are simpler than the ones given in [18].
This leads in general, also at other places where recently other aspects of compact and discrete quantum groups are studied, to simplifications and a better understanding. This should not come as a surprise because it is already the case in the approach to discrete quantum groups itself as treated in [18].
@@
Taken abstraction of this (and of the historical context), people who want to work with compact and discrete quantum groups can consider any of the two approaches. To a great extend, it is a matter of taste which you take as a starting point (although we are convinced that starting with discrete quantum groups is a bit easier). However, for a full understanding, we feel it is worthwhile to study the two theories also as independent entities.